OpenStudy (anonymous):

find the local extrema, the intervals on which the function is increasing and the intervals on which the function is decreasing g(x)=x^2-x-12

OpenStudy (aum):

Find g'(x). If g'(x) > 0, the function is increasing. If g'(x) < 0, the function is decreasing.

OpenStudy (aum):

solve g'(x) = 0 to find the local extrema.

OpenStudy (anonymous):

so g'(x)=2x-1 set that equal to 0?

OpenStudy (anonymous):

so one critical point is 1/2

OpenStudy (anonymous):

oh wait that would be a local max right?

OpenStudy (aum):

Correct. x = 1/2 is a critical point. The function is a parabola with a positive leading coefficient. Therefore, the parabola opens upward. So x = 1/2 is a local minima. g'(x) > 0 gives the interval where the function is increasing. Solve 2x - 1 > 0

OpenStudy (anonymous):

so x > 1/2 which is increasing

OpenStudy (aum):

Yes. And and for x < 1/2 the function is decreasing.

OpenStudy (anonymous):

so the interval would be \[(-\infty, 1/2)U(1/2, \infty)\]

OpenStudy (anonymous):

wait.. no it there wouldnt be a union

OpenStudy (anonymous):

so decreasing is \[(-\infty, 1/2)\] and increasing at \[(1/2, \infty)\]

OpenStudy (aum):

Correct.

OpenStudy (anonymous):

cool thanks

OpenStudy (aum):

Local extrema at x = 1/2. It is a minima and the minimum value is f(1/2) = (1/2)^2 - 1/2 - 12 = -12.25

OpenStudy (aum):

you are welcome.